# From Grounding to Manipulation: Case Studies of Foundation Model Integration in Embodied Robotic Systems

### v1

@@ Abstract
Foundation models (FMs) are increasingly used to bridge language and action in embodied agents, yet the operational characteristics of different FM integration strategies remain under-explored—particularly for complex instruction following and versatile action generation in changing environments. This paper examines three paradigms for building robotic systems: end-to-end vision-language-action (VLA) models that implicitly integrate perception and planning, and modular pipelines incorporating either vision-language models (VLMs) or multimodal large language models (LLMs). We evaluate these paradigms through two focused case studies: a complex instruction grounding task assessing fine-grained instruction understanding and cross-modal disambiguation, and an object manipulation task targeting skill transfer via VLA finetuning. Our experiments in zero-shot and few-shot settings reveal trade-offs in generalization and data efficiency. By exploring performance limits, we distill design implications for developing language-driven physical agents and outline emerging challenges and opportunities for FM-powered robotics in real-world conditions.
@@
Xiuchao Sui1thanks: xiuchao.sui@gmail.com.  and Daiying Tian1thanks: Equal contribution. Corresponding author.  and Qi Sun 2  and Ruirui Chen1  and Dongkyu Choi 1and Kenneth Kwok 1and Soujanya Poria 2 1Institute of High Performance Computing, A*STAR, Singapore 2Singapore University of Technology and Design, Singapore
@@ 1 Introduction
Natural language is emerging as a universal interface for embodied robotic systems. Recent foundation models (FMs) allow robots to follow free-form instructions in perception, reasoning, and motor commands, offering the promise of language-grounded autonomy. They include multimodal large language models (LLMs) Grattafiori et al. (2024); Bai et al. (2025); Lu et al. (2024), vision–language models (VLMs) Liu et al. (2024a); Ravi et al. (2025); Ren et al. (2024); Li et al. (2023), and vision–language–action (VLA) models Kim et al. (2024); Zheng et al. (2025); Qu et al. (2025); Bu et al. (2025).
@@
However, turning the promise of language-grounded autonomy into deployable systems is highly challenging. Robots must (i) map ambiguous instructions to the physical world (instruction grounding), (ii) execute reliably across novel objects, scenes, and robot morphologies (generalizable execution), and (iii) achieving the aforementioned goals with limited data (efficient adaptation). How well different FM integration strategies meet these competing requirements remains under-explored (Fig. 1).
Figure 1: Challenges of foundation models in embodied robotic systems include cross-modal instruction grounding, generalization across environments and morphologies, and data-efficient adaptation for the real world.
@@
To our best knowledge, this work delivers the first head-to-head empirical comparison of three prevalent integration paradigms: end-to-end VLAs that directly map language and vision to actions, multimodal LLM agents that orchestrate perception and control through tool calls, and modular VLM pipelines that couple perception-specialist FMs with task-specific planners (Fig. 2; Table 1).
@@
We evaluate these paradigms through tabletop case studies designed to highlight their complementary strengths and limitations. Specifically, we consider two task categories: (i) Complex Instruction Grounding, which probes fine-grained understanding and cross-modal disambiguation (Sec.3); and (ii) Object Manipulation, which measures the ability to transfer learned skills after VLA fine-tuning under distribution shifts, complemented by comparative and ablation studies (Sec.4).
Our zero-shot grounding experiments reveal distinct trade-offs across integration strategies. VLM pipelines prioritize interpretability and data efficiency, sacrificing flexibility and peak performance. While underperforming on handling complex instruction grounding, they deliver moderate performance in object grounding—using less than 1% of the parameters required by multimodal LLMs. In contrast, multimodal LLM agents generalize better on complex instructions but incur significantly higher inference costs. Notably, smaller reasoning-focused models such as GPT-4o-mini can even outperform larger models like GPT-4o on certain tasks. VLAs, with their tightly coupled perception-to-action pathways, support streamlined action generation, yet struggle to reason about rare or abstract concepts. We further examine the trade-off between model size and performance by analyzing quantization effects on open-source multimodal LLMs. These findings offer practical guidance for developing language-driven robotic systems under real-world constraints.
Within the VLA paradigm, we categorize models by their action generation mechanisms—autoregressive Vaswani et al. (2017); Kim et al. (2024); Hung et al. (2025) and diffusion-based approaches Ho et al. (2020); Chi et al. (2023); Reuss et al. (2024); Wen et al. (2024). We evaluate their adaptability through fine-tuning under distribution shifts, mirroring real-world deployment scenarios. To assess generalization, we analyze robustness to environmental perturbations and variation in object appearance and robot morphology.
To summarize, the main contributions of this work are as follows:
  • To our best knowledge, we present the first systematic comparison of end-to-end VLA, modular VLM, and modular multimodal LLM architectures on the same set of embodied tasks.
@@
  • We release a dataset and accompanying code [^1] https://github.com/xiuchao/InstructionGrounding that supports the evaluation of complex instruction grounding and manipulation transfer—covering accordance comprehension, cross-modal reasoning, and motor adaptation.
  • We benchmark state-of-the-art VLAs and multimodal LLMs, offering timely insight into the capabilities and failure modes of modern FMs in robotics.
  • We distill actionable trade-offs that practitioners can apply when choosing an FM stack for language-driven embodied agents.
@@
  • We release a complete, end-to-end claw-machine robot system to demonstrate FM integrations in real-world applications [^2] https://github.com/HRItdy/claw_machine.
@@ 2 Foundation Model Integration for Language-Guided Robotics
Figure 2: Foundation model integration strategies in language driven robots: (a) End-to-end VLA models, (b) Modular VLM pipelines, and (c) Multimodal LLM agents. Each strategy reflects a distinct interface between language, perception, and control.
2.1 End-to-End Vision-Language-Action
@@
Vision-Language-Action (VLA) models operate in an end-to-end manner, directly translating visual observations and natural language instructions into low-level actions without decoupled perception, language, and control modules (Fig. 2a). Two mainstream paradigms have emerged within this framework: auto-regressive and diffusion-based action generation. Through large-scale pretraining, these models acquire broad capabilities that support generalization across tasks. However, efficient adaptation to real-world settings remains a significant challenge.
Autoregressive VLA Models.
Autoregressive VLA models typically process language and visual inputs, employing various tokenization strategies to convert multimodal data into a unified latent space. Then the transformer-based decoder generates actions step-by-step in an autoregressive manner conditioned on the input context and previously generated actions, allowing structured action generation and planning.
Building on the previous groundbreaking models such as RT-1 Brohan et al. (2023), RT-2 Zitkovich et al. (2023) and VIMA Jiang et al. (2023), OpenVLA Kim et al. (2024) emerges as an important open-source method, combining a fine-tuned Llama 2 (7B) model with DinoV2 and SigLIP for visual tokenization, pretrained on the Open-X-Embodiment dataset Collaboration et al. (2024) consisting of 970k real-world robot demonstrations. TraceVLA Zheng et al. (2025) improves OpenVLA with visual trace prompting to enhance spatiotemporal awareness. Emma-X Sun et al. (2024) further refines dataset quality using a trajectory segmentation and Embodied Chain-of-Thought reasoning. This work is followed by NORA Hung et al. (2025), which uses Qwen-2.5-VL-3B as backbone and pretrained on the Open-X-Embodiment dataset. In parallel, other efforts integrate tactile sensing into robot perception Yang et al. (2024); Zhao et al. (2024), supporting more generic robot policies. Other recent developments include distilling spatial representations from VLMs, e.g., Gemini-Robitics Gemini Robotics Team (2025), and enriching training with additional features, such as 3D spatial relationships in SpatialVLA Qu et al. (2025), task-centric latent space in UniVLA Bu et al. (2025), and integrating multimodal understanding with action prediction in ChatVLA Zhou et al. (2025) and UP-VLA Zhang et al. (2025).
Diffusion-based VLA Models.
Diffusion-based VLA models formulate action generation as a denoising process over latent trajectories. Given a noisy version of a full action sequence, the model learns to recover the true trajectory conditioned on language and vision.
Diffusion Policy (DP) Chi et al. (2023) pioneered the use of diffusion model for visuomotor policy representation, laying the foundation for subsequent multimodal approaches. Octo Team et al. (2024) uses conditional diffusion decoding for action sequence prediction. Rather than lightweight diffusion heads, Reuss et al. (2024); Wen et al. (2024); Li et al. (2024b) use larger and more dedicated diffusion policy modules the as action decoder. DTP Fan et al. (2025) introduces a trajectory-level guidance to enhance diffusion-based planning. Recent works such as $\pi_{0}$ Black et al. (2024) and $\pi_{0.5}$ Intelligence et al. (2025) integrate a pretrained VLM with a flow-matching-based action expert to model action distributions. These models often utilize FAST Pertsch et al. (2025) for efficient compressed action tokenization. HybridVLA Liu et al. (2025) unifies diffusion and autoregressive action prediction within a single LLM-based framework. A growing focus is placed on adapting VLA models to diverse embodiments including bimanual manipulation and humanoid robots, such as RDT-1b Liu et al. (2024b), DexVLA Wen et al. (2025) and GR00T N1 Bjorck et al. (2025).
@@
VLA models provide a unified, end-to-end framework for robotic manipulation, (i) seamlessly integrating visual, language, and action modalities. Leveraging large-scale pretraining and fine-tuning, these models exhibit (ii) strong potential for generalizing across diverse manipulation tasks and robotic embodiments. However, their performance is constrained by the (iii) limited availability of high-quality, diverse robotic datasets. Pretraining can also introduce biases from training distributions, leading to (iv) degraded performance in out-of-distribution scenarios, such as novel tasks or different robotic embodiments. Therefore, despite their potential, further work is needed to enhance their robustness and generalization across real-world settings, for example, efficient adaptation using few-shot data.
2.2 Modular Vision–Language Pipelines
@@
In a modular vision-language pipeline (Fig. 2b), perception is handled by a specialist vision language model (VLM) that outputs symbolic scene information, typically grounded 2-D / 3-D bounding boxes, segmentation masks, or referring expression pointers. A downstream planner or policy module then consumes this structured representation to generate low-level actions. The language channel is therefore disentangled from motor control, allowing each module to be tuned independently, thus preserving the transparency and plug-and-play advantages of classical planning.
@@
Modular VLM pipelines strike a balance between transparency and adaptability, and delivers practical benefits: (i) interpretability—detections can be inspected; (ii) lightweight—the model parameters are usually around 100Msimilar-$\sim$600M, approximately 1% $\sim$ 6% the size of LLaMA 3.2 Vision 11B Grattafiori et al. (2024). On the other hand, it is limited at (i) interaction rigidness compared with more flexible multimodal LLMs, and (ii) pipeline brittleness where perception errors propagate without mitigation (Fig. 2b; Table 1). Their success hinges on robust open-vocabulary grounding—precisely the capability our Instruction Grounding case study stresses in Section 3.
@@ 2.3 Multimodal LLM Agents as Orchestrators
Multimodal LLM agents place a large, tool-calling language model at the centre of the control loop (Fig. 2c). The LLM receives raw user utterances, selectively invokes vision tools (e.g., a detector or depth estimator) via function calls, reasons over their outputs in-context, and finally issues high-level action primitives to a low-level controller. The agent therefore acts as a cognitive hub that binds perception and control through natural language.
@@
Multimodal LLMs are taking increasingly important roles in robotics. Gemini Robotics Gemini Robotics Team (2025) integrates perception, spatial reasoning, and trajectory synthesis into one Gemini-2.0 backbone Google DeepMind (2024), which serves as the embodied brain. Li et al. (2024c), in the same vibe as Gemini Robotics, leverages the inherent common sense and reasoning capabilities of these models by fine-tuning adapter modules through a chain-of-thought training paradigm. It endows the model with accurate pose prediction and precise manipulation abilities. These works collectively show the trend that the multimodal LLM shifts to the “cognitive hub” in robot systems. Glocker et al. (2025) build a modular agent-orchestration system for household object management robots. It utilize Llama 3.2 Vision Grattafiori et al. (2024) for open-vocabulary perception to facilitate creating grounded task plans, while the limitations of the multmodal LLM were not discussed.
Somewhat similar to our work, Li et al. (2024a) investigates the eligibility of Multimodal LLMs to serve as the “brain” for in-home robotics by providing a benchmark to compare models along the axes of perception, visual reasoning and task planing. Models like GPT-4V, Qwen-VL Bai et al. (2025) and DeepSeek-VL Lu et al. (2024) were included, but more recent releases were not covered—likely due to the fact that the field is moving fast with new models emerging in rapid succession.
@@
Multimodal LLM agents excel in (i) visual commonsense reasoning, leveraging extensive language priors to generalize to novel concepts beyond the reach of most specialist VLMs, and (ii) instruction following with support for fine-grained visual understanding and dynamic planning. Despite their expressive power, however, these models are (iii)  resource-intensive, posing challenges for deployment—particularly on mobile robotic platforms.
@@ 3 Case Studies on Instruction Grounding
Natural language instruction grounding involves translating user intents into clear, actionable goals in a visual scene, which is a key capability for embodied AI Gemini Robotics Team (2025). Our case study offers empirical insights into the grounding performance of various models through the lens of challenging cross-modal disambiguation, and further examines the trade-offs introduced by model sizes and quantization—providing practical suggestions for efficient deployment.
@@
To minimize the impact of vision priors on measured performance, we design benchmarking scenarios using household objects placed on a tabletop. These objects are commonly represented in the training datasets of the foundation models, and the tabletop setup features minimal variation in lighting and camera angles—ensuring that the evaluation primarily reflects grounding capabilities.
We curated a new Instruction Grounding benchmark dataset. In images containing multiple household objects, each object is tagged with a number as the visual prompt, and each image is paired with language instructions crafted to test visual commonsense and cross-modal disambiguation concerning attribute or spatial relationships – for “pick up the red-capped marker,” the color must be used to select one among a few markers; whereas “grasp the cup in front of the screwdriver” requires reasoning over spatial relations (Fig. 3; Appendix).
Language ambiguity often leads to execution failure in an embodied system. By comparing specialist VLMs and multimodal LLMs, we reveal their concrete failure modes that further inform our design implications in Sec. 5.
Model Easy Medium Hard Avg Specialist VLMs GroundingDINO-86M 0.518 0.357 0.349 0.408 GroundingDINO-145M 0.443 0.320 0.355 0.372 Closed-source Multimodal LLMs Gemini2.5-Pro-Exp 0.904 0.765 0.793 0.821 Gemini2.0-Flash 0.884 0.738 0.678 0.767 GPT-4.5 0.837 0.723 0.739 0.766 GPT-4o 0.814 0.745 0.683 0.747 GPT-4o-mini 0.803 0.722 0.604 0.710 o4-mini 0.721 0.769 0.710 0.733 GPT-4V 0.470 0.476 0.467 0.471 Open-source Multimodal LLMs Llama-3.2V-90B 0.722 0.701 0.657 0.693 Llama-3.2V-11B 0.583 0.569 0.547 0.566 Llama-4-Maverick 0.698 0.576 0.634 0.636 Llama-4-Scout 0.776 0.615 0.624 0.672 Qwen2-VL-72B 0.686 0.614 0.558 0.619 Gemma-3-27B 0.452 0.384 0.267 0.368 DS-Janus-Pro-7B 0.444 0.330 0.317 0.364 Phi-3.5-Vision-4.2B 0.291 0.357 0.205 0.284
Table 2: Object grounding performance of specialist VLMs and multimodal LLMs (closed-source and open-source) across varying scene complexity levels. Models are evaluated on easy, medium, and hard cluttered scenes, with macro accuracy reported.
Figure 4: Performance of complex instruction grounding across VLM–LLM pipelines and end-to-end multimodal LLMs. Macro accuracy is reported across instruction types—implicit, attribute-based, and relationship-based. Subfigures show (a) proprietary models and (b) open-source models along with their Int4-quantized variants.
@@
We begin with a foundational question for instruction grounding: Can FMs accurately recognize objects in cluttered open scenes? Table 2 presents the performance hierarchy for specialist VLMs and a range of multimodal LLMs, serving as a basis for deeper analysis of ambiguity resolution in later sections.
  • Despite their popularity in modular pipelines, GroundingDINO achieve only 35–41% accuracy, as it struggles with featureless objects, e.g. ‘the can’. Moreover, it is brittle in open scenes, e.g. a ‘screwdriver’ is constantly recognized as a ‘marker’, which instead is an easy case for multimodal LLMs which embodied large volume of visual commonsense.
  • Gemini 2.5-Pro takes the first place with 0.82, followed by Gemini 2.0-Flash and GPT-4.5. Open-source systems still trail the proprietary tier. Llama 3.2-Vision 90B reaches about 84% of Gemini 2.5’s score, while the more recent Llama 4 releases did not outperform it.
  • Smaller community models (Gemma-27B, Phi-Vision) fall below the specialist-threshold, suggesting that they are still inadequate for fine-grained grounding in cluttered scenarios.
  • Last, when compute is a bottleneck, GPT-4o-mini (0.71) and Llama 3.2-Vision 11B (0.57) provide the best speed–accuracy trade-off, delivering decent performance without incurring heavy memory footprint or high API costs.
@@
This task is framed as a multiple-choice problem, where the model is asked to select the correct object index in a cluttered scene based on three types of natural language instructions: implicit, attribute-based, and relationship-based—each type probes a distinct grounding challenge. We evaluate a series of multimodal LLMs, using a modular VLM–LLM pipeline as a baseline. In this pipeline, the LLM parses the instruction to infer likely targets, queries GroundingDINO to detect candidate objects, and selects from the detected boxes—essentially guessing without directly perceiving the scene.
  • Implicit Instruction Grounding. Instructions like “I need a tool to tighten the screws” only refer to the target object implicitly, and the model needs to infer the target object using its common sense priors. For such instructions, the modular VLM–LLM pipeline struggles to select a screwdriver, lacking embedded affordance reasoning. In contrast, multimodal LLMs perform well, reflecting strong visual commonsense. GPT-4.5 demonstrates exceptional performance (0.94), though its high inference cost—20×20$\times$20 × that of Gemini 2.5 makes it cost-prohibitive for most applications.
  • Relational Reasoning Remains Challenging. This category requires resolving referential ambiguity through implicit chain-of-thought reasoning: grounding objects, modeling spatial relationships, and disambiguating targets (e.g., identifying the correct mug among many based on “next to something”). Accuracy drops significantly nearly across all models. Only Gemini 2.5-Pro and o4-mini achieve accuracy above 0.80—the former likely benefits from embodied training data, while the latter demonstrates strong reasoning capabilities. Notably, o4-mini is a medium-sized model, yet it outperforms larger models like GPT-4.5 on relational instructions—suggesting that structured reasoning may help close, or even overcome the performance gap brought by different model scales.

### v2

@@ Abstract
Foundation models (FMs) are increasingly applied to bridge language and action in embodied agents, yet the operational characteristics of different integration strategies remain under-explored—especially for complex instruction following and versatile action generation in changing environments. We investigate three paradigms for robotic systems: end-to-end vision-language-action models (VLAs) that implicitly unify perception and planning, and modular pipelines using either vision-language models (VLMs) or multimodal large language models (MLLMs). Two case studies frame the comparison: instruction grounding, which probs fine-grained language understanding and cross-modal disambiguation; and object manipulation, which targets skill transfer via VLA finetuning. Our experiments reveal trade-offs in system scale, generalization and data efficiency. These findings indicate design lessons for language-driven physical agents and point to challenges and opportunities for FM-powered robotics in real-world conditions.
@@
Xiuchao Sui1thanks: Equal contribution. Corresponding authors. {Sui_Xiuchao, Tian_Daiying}@a-star.edu.sg  and Daiying Tian [^1] footnotemark: 1  and Qi Sun 2  and Ruirui Chen1  and Dongkyu Choi 1and Kenneth Kwok 1and Soujanya Poria 2 1IHPC, Agency for Science, Technology and Research, Singapore 2Nanyang Technological University, Singapore
@@ 1 Introduction
Natural language is emerging as a universal interface for embodied robotics. Advances in foundation models (FMs) enable robots to follow free-form instructions across perception, reasoning, and motor control, offering the promise of language-grounded autonomy. These models include vision–language models (VLMs) Liu et al. (2024a); Ravi et al. (2025); Ren et al. (2024); Li et al. (2023), multimodal large language models (MLLMs) Grattafiori et al. (2024); Bai et al. (2025); Lu et al. (2024), and vision–language–action (VLA) models Kim et al. (2024); Zheng et al. (2025); Qu et al. (2025); Bu et al. (2025).
@@
However, realizing the promise of language-grounded autonomy in deployable systems remains highly challenging. Robots must (i) map ambiguous instructions to the physical world (instruction grounding), (ii) execute reliably across novel objects, scenes, and morphologies (generalizable execution), and (iii) achieve these goals with limited data (efficient adaptation). How effectively different FM integration strategies address these competing requirements remains under-explored (Fig. 1).
Figure 1: Key challenges of FMs in embodied robotics.
@@
This work offers an empirical study of three integration paradigms: end-to-end VLAs mapping language and vision to actions, MLLM agents orchestrating perception and control, and modular VLM pipelines pairing perception-specialist FMs with task-specific planners (Fig.2; Table1).
@@
We evaluate these paradigms through tabletop case studies that expose their complementary strengths and limitations. Two task categories are considered: complex instruction grounding, probing fine-grained understanding and cross-modal disambiguation (Sec.3); and object manipulation, assessing skill transfer after VLA fine-tuning under distribution shifts, complemented by comparative and ablation analyses (Sec.4).
Our instruction grounding experiments reveal distinct trade-offs across integration strategies. VLM pipelines emphasize interpretability and efficiency, but fall short of peak performance. They struggle with complex instruction grounding yet achieve moderate object grounding—using less than 1% of the parameters required by MLLMs. In contrast, MLLMs generalize better on complex instructions but incur substantially higher inference costs, with smaller reasoning-focused models at times outperforming larger ones. VLAs, with tightly coupled perception-to-action pathways, streamline action generation but make it difficult to isolate and assess grounding performance. To probe this gap, we evaluate their perception heads as reference points, which also surfaces open questions for VLA architecture design.
While MLLMs perform well on grounding tasks, their scale limits practical deployment on robotic platforms. Quantization is often adopted as a simple remedy, yet instruction-dependent behaviors emerge, with certain reasoning capabilities degrading disproportionately under naïve compression. This underscores the need for fine-grained quantization strategies. Taken together, our findings clarify the trade-offs between model scale and performance and provide practical guidance for building robotic systems under real-world constraints.
Our object manipulation experiments investigate the skill-adaptation capabilities of VLAs in both real-world and simulated settings. We examine how these models transfer manipulation skills under distribution shifts, evaluating their training dynamics, generalization, and robustness. Results reveal fragile training and slow adaptation in generalist policies, such as OpenVLA Kim et al. (2024), compared to task-specific counterparts, underscoring the challenges of real-world deployment.”
In summary, our main contributions are:
  • We analyse three FM integration paradigms on shared embodied tasks designed to probe the capabilities and trade-offs of FMs.
@@
  • We release a dataset and code [^1] https://github.com/xiuchao/InstructionGrounding for evaluating instruction grounding and object manipulation, covering cross-modal reasoning and skill adaptation under varied layouts.
  • We providing timely insights into state-of-the-art VLAs and MLLMs, investigating their capabilities and failure modes and distilling practical trade-offs to guide practitioners in selecting FM stacks for language-driven embodied agents.
@@
  • We also release a complete end-to-end claw-machine robot system as a real-world FM integration demo [^2] https://github.com/HRItdy/claw_machine.
@@ 2 Foundation Model Integration for Language-Guided Robotics
Figure 2: Three FM integration strategies for embodied robotics, highlighting distinct interfaces between language, perception, and control.
2.1 End-to-End VLA Models
@@
VLAs operate in an end-to-end manner, directly translating visual observations and natural language instructions into low-level actions without decoupled perception, language, and control modules (Fig. 2a). Two mainstream paradigms have emerged within this framework: auto-regressive and diffusion-based action generation. Through large-scale pretraining, these models acquire broad capabilities that support generalization across tasks. However, efficient adaptation to real-world settings remains a significant challenge.
Autoregressive Models.
These models generate actions step by step, with each action conditioned on the current perceptual input and previously generated outputs. These models typically encode visual observations and language into a shared latent space, then employ a transformer-based decoder to autoregressively predict low-level control tokens, such as joint angles, end-effector displacements and gripper states, which are typically obtained by discretizing continuous motor signals into token sequences.
RT-1 Brohan et al. (2023) and RT-2 Zitkovich et al. (2023) demonstrated the effectiveness of large-scale pretraining for task generalization. Building on these scaling successes, OpenVLA Kim et al. (2024) became a landmark open-source effort, combining Llama-2 with DinoV2 and SigLIP, trained on the large-scale Open-X-Embodiment dataset Collaboration et al. (2024). Subsequent works such as Emma-X Sun et al. (2024), NORA Hung et al. (2025) and TraceVLA Zheng et al. (2025), further advanced this line through dataset quality, stronger backbones and improved spatiotemporal prompting.
Subsequent models further incorporate richer modalities such as tactile sensing Yang et al. (2024); Zhao et al. (2024), structural priors including spatial representations distilled from VLMs Gemini Robotics Team (2025) and 3D spatial relationships in SpatialVLA Qu et al. (2025), and refinements to the embodied latent space, as in UniVLA Bu et al. (2025), which learns task-specific latent representations.
Diffusion-based Models.
They generate actions by progressively denoising trajectories from noise, modeling the distribution of future actions as a whole rather than step by step. Like autoregressive VLAs, these models encode visual observations and language into a latent representation, but instead apply a diffusion process that iteratively refines action sequences into feasible trajectories. Diffusion-based models trade inference speed for greater trajectory coherence.
Diffusion Policy Chi et al. (2023) pioneered the use of diffusion models for visuomotor policy learning, showing improved global consistency and robustness compared to autoregressive baselines. Octo Team et al. (2024) applies conditional diffusion decoding for action sequence prediction. Subsequent models further scales diffusion heads into larger, dedicated policy modules Reuss et al. (2024); Wen et al. (2024); Li et al. (2024b), and integrate trajectory-level guidance to improve long-horizon planning Fan et al. (2025). Notable models like $\pi_{0}$ Black et al. (2024) combine pretrained MLLM (PaliGemma-3B) with flow-matching-based action experts to produce continuous, high-frequency control across robot embodiments. $\pi_{0.5}$ Intelligence et al. (2025) builds on this by co-training across diverse environments and data modalities to improve generalization to new settings.
Beyond architectural advances, diffusion-based models increasingly targets diverse embodiments, extending from single-arm manipulation to bimanual systems and humanoid robots. Representative examples include RDT-1B Liu et al. (2024b), DexVLA Wen et al. (2025), and GR00T N1 Bjorck et al. (2025), which demonstrate the scalability of FM-driven action generation to more complex morphologies.
@@
Leveraging large-scale pretraining, VLAs hold the potential to generalize across diverse manipulation tasks and robot morphologies. Yet progress is constrained by the scarcity of high-quality, diverse robotic datasets, which limits both scale and coverage. Pretraining can also propagate biases from training distributions, causing degraded performance on novel tasks, in unseen environments, or with unfamiliar embodiments. Despite their promise, VLAs remain brittle in real-world deployment, as highlighted in our Object Manipulation case study (Section 4), underscoring the need for better data curation, bias mitigation, and efficient adaptation strategies such as few-shot and continual learning.
2.2 Modular VLM Pipelines
@@
In this paradigm (Fig. 2b), perception is handled by a specialist VLM that outputs symbolic scene information, typically grounded 2D/3D bounding boxes, segmentation masks, or referring expression pointers. A downstream planner or policy module then consumes this structured representation to generate low-level actions. The language channel is therefore disentangled from motor control, allowing each module to be tuned independently, thus preserving the transparency and plug-and-play advantages of classical planning.
@@
[…]
@@ 2.3 Multimodal LLM Agents as Orchestrators
In this paradigm (Fig. 2c), MLLMs take raw user utterances, selectively invoke vision tools (e.g., a detector or depth estimator) via function calls, reason over their outputs in context, and issue high-level action primitives to a low-level controller. An MLLM agent thus places a large tool-calling language model at the center of the control loop, acting as a cognitive hub that binds perception and control through natural language.
@@
MLLMs are playing increasingly important roles in robotics. Gemini Robotics Gemini Robotics Team (2025) integrates perception, spatial reasoning, and trajectory synthesis into a single Gemini-2.0 backbone Google DeepMind (2024), serving as an embodied brain. ManiLLM Li et al. (2024c), in a similar spirit, leverages the common-sense and reasoning capabilities of MLLMs by fine-tuning adapter modules with a chain-of-thought training paradigm, enabling accurate pose prediction and precise manipulation. These works illustrate the emerging trend of MLLMs shifting toward the role of a cognitive hub in robot systems. Hub-LLaMA Glocker et al. (2025) further builds a modular agent-orchestration system for household object management, using LLaMA 3.2 Vision Grattafiori et al. (2024) for open-vocabulary perception to ground task plans, though its limitations are not discussed.
Somewhat related to our work, Li et al. (2024a) evaluates the suitability of MLLMs as a “brain” for in-home robotics, providing a benchmark that compares models across perception, visual reasoning, and task planning. The benchmark included a few MLLMs available at the time, while newer releases were not covered—illustrating the rapid pace of progress in this area.
@@
MLLMs excel in (i) visual commonsense reasoning, leveraging extensive language priors to generalize to novel concepts beyond the reach of most specialist VLMs, and (ii) instruction following with support for fine-grained visual understanding and dynamic planning. Despite their expressive power, however, these models are (iii)  resource-intensive, posing challenges for deployment—particularly on mobile robotic platforms. We further examine the capability limits and trade-offs of MLLMs (Section 3).
@@ 3 Case Studies on Instruction Grounding
Natural language instruction grounding translates user intent into actionable goals within a visual scene. Our case study evaluates grounding performance under challenging cross-modal disambiguation, highlighting trade-offs and offering guidance for efficient deployment.
@@
To isolate grounding ability from general vision priors, we design carefully controlled scenarios using household objects placed on a tabletop. These objects are widely represented in FM training corpora, while the tabletop setup minimizes variation in lighting and camera angles—ensuring the evaluation primarily reflects grounding performance.
We curated a new Instruction Grounding benchmark (Fig. 3) consisting of images with multiple household objects, each tagged with numbers as visual prompts and paired with instructions targeting visual commonsense and cross-modal disambiguation. For example, “pick up the red-capped marker” requires attribute reasoning to choose among markers, while “grasp the cup in front of the screwdriver” tests spatial reasoning (Appendix Table 4). These tasks highlight how grounding errors can lead to execution failures in embodied systems.
Model Easy Medium Hard Avg Specialist VLMs GroundingDINO-86M 0.518 0.357 0.349 0.408 GroundingDINO-145M 0.443 0.320 0.355 0.372 Proprietary MLLMs Gemini2.5-Pro-Exp 0.904 0.765 0.793 0.821 Gemini2.0-Flash 0.884 0.738 0.678 0.767 GPT-5-auto 0.881 0.829 0.760 0.823 GPT-5-mini 0.749 0.737 0.776 0.754 GPT-4o 0.814 0.745 0.683 0.747 GPT-4o-mini 0.803 0.722 0.604 0.710 o4-mini 0.721 0.769 0.710 0.733 Open-source MLLMs Llama-3.2V-90B 0.722 0.701 0.657 0.693 Llama-3.2V-11B 0.583 0.569 0.547 0.566 Llama-4-Maverick 0.698 0.576 0.634 0.636 Llama-4-Scout 0.776 0.615 0.624 0.672 Qwen2-VL-72B 0.686 0.614 0.558 0.619 Gemma-3-27B 0.452 0.384 0.267 0.368 DS-Janus-Pro-7B 0.444 0.330 0.317 0.364 Phi-3.5-Vision-4.2B 0.291 0.357 0.205 0.284 Base MLLMs in VLAs PaliGemma-3B ($\pi_{0}$) 0.118 0.07 0.05 0.079 QwenVL-3B (NORA) 0.519 0.535 0.577 0.543
Table 2: Object grounding performance of specialist VLMs and MLLMs across cluttered scenes of varying complexities, with macro accuracy reported.
Figure 4: Performance of complex instruction grounding across modular VLM pipeline and MLLMs. Macro accuracy is reported across instruction types—implicit, attribute-based, and relationship-based. Subfigures show (a) proprietary models and (b) open-source models along with their Int4-quantized variants.
@@
We start with a necessary first question for instruction grounding: Can FMs reliably recognize objects in cluttered, open-world scenes? Table 2 highligh key observations and point to opportunities for VLA design.
  • Despite their popularity in modular pipelines, GroundingDINO achieves around 0.3-0.4 accuracy, as it struggles with featureless objects, e.g. ‘the can’. Moreover, it is brittle in open scenes, e.g. a ‘screwdriver’ is constantly recognized as a ‘marker’, which instead is an easy case for MLLMs which embodied large volume of visual commonsense (Appendix Fig. 10).
  • Gemini 2.5-Pro and GPT-5 achieve top performance with an average score of 0.82, followed by Gemini 2.0-Flash and GPT-4.5. Open-source models continue to lag behind, with LLaMA 3.2-Vision 90B reaching 84% of the performance of top proprietay models.
  • Small- and medium-scale models (e.g., Gemma-27B, Phi-Vision) generally fall below the specialist threshold, underscoring their limitations for fine-grained object grounding in open scenes. Notable exceptions include GPT5-mini (0.754) and QwenVL-3B (0.543), which offer favorable speed–accuracy trade-offs, achieving reasonable performance without high API costs or heavy memory footprint of larger models.
  • PaliGemma-3B in $\pi_{0}$ struggles to follow structured prompts and can only follow simpler instructions. Its grounding capability is highly limited, often detecting just a single object per scene. Reported accuracy is therefore approximate, based on a looser evaluation criterion. While QwenVL-3B adopted in NORA, reaches higher grounding accuracy, which may partially explain its stronger performance on out-of-distribution manipulation tasks (Appendix Table 6).
  • Two design questions emerges: (i) how should perception modules be selected or adapted to improve instruction-following in embodied systems? and (ii) can compact, grounding-capable perception modules be distilled from large-scale models to support efficient deployment?
@@
This task is framed as a multiple-choice problem, where the model selects the correct object index in a cluttered scene given three types of natural language instructions: implicit, attribute-based, and relationship-based. Each type probes a distinct grounding challenge (Fig. 4). As a baseline, we evaluate a modular VLM pipeline in which GPT-4 parses the instruction to infer likely targets, queries GroundingDINO to detect candidate objects, and selects from the detected boxes—essentially guessing without directly perceiving the scene.
[…]
  • Relational Reasoning Remains Challenging. This category requires resolving referential ambiguity through implicit chain-of-thought reasoning: grounding objects, modeling spatial relationships, and disambiguating targets (e.g., identifying the correct mug among many based on “next to something”). Accuracy drops significantly nearly across all models. GPT-5, Gemini 2.5-Pro and o4-mini achieve accuracy above 0.80—demonstrating the benefits from embodied training data and strong reasoning capabilities. Notably, o4-mini is a medium-sized model, yet it outperforms larger models like GPT-4o on relational instructions—suggesting that structured reasoning may help close, or even overcome the performance gap brought by different model scales.